At the beginning of 2018, Lindy Weink, the controller of Sunland Inc., reviewed the expected useful life and residual value of two of the companys machines and proposed changes as follows Useful Life (in years) Residual Value Machine Date AcquiredCost Original Proposed Original Proposed $42,520 $58,530 $5,110 $3,332 #1 #2 Jan. 1, 2008 Jan. 1, 2016 $752,700 $126,490 20 5 25 Calculate the annual depreciation for each asset using the straight-line method and the original useful life and residual value. Machine 1 Machine 2 LINK TO TEXT Calculate the accumulated depreclation and carrying amount of each asset on December 31, 2017 Machine1 Machine 2 Accumulated depreciation Carrying amount
Answer #1

Calculate annual depreciation :

Machine 1 = (752700-42520/20) = 35509

Machine 2 = (126490-5110/5) = 24276

Calculate accumlated depreciation and carrying amount of each asset on december 31,2017

Machin 1 Machine 2
Accumlated depreciation 35509*10 = 355090 24276*2=48552
Carrying amount (752700-355090) = 397610 (126490-48552) = 77938